OpenStudy (anonymous):

Solve for x: 3x/4 =x+2/3

OpenStudy (pottersheep):

try multiplying everything by 12

OpenStudy (pottersheep):

[3x/4 =x+2/3] 12 9x = 12x + 8 See, we got rid of the fractions

OpenStudy (pottersheep):

9x = 12x + 8 -3x = 8 x=-8/3

OpenStudy (anonymous):

Oh I see! Thank you so much!!

OpenStudy (pottersheep):

Glad I could help